Five bulbs,each rated as 220v 100 watt burnt for 4 hours.What is the cost of electricity consumed if the rate is 50 paise per unit?
1
Expert's answer
2012-07-26T08:37:56-0400
The power consumed is

W = 5 * 100 watt * 4 hours = 5 * 0.100 kilowatt * 4 hours = 2 kilowatt*hours.

So, the cost of consumed electricity is

P = 2 units * 50 paise = 1 rupee.